### v1

# Adaptive debiased SGD in high-dimensional GLMs with steaming data

Abstract
Online statistical inference facilitates real-time analysis of sequentially collected data, making it different from traditional methods that rely on static datasets. This paper introduces a novel approach to online inference in high-dimensional generalized linear models, where we update regression coefficient estimates and their standard errors upon each new data arrival. In contrast to existing methods that either require full dataset access or large-dimensional summary statistics storage, our method operates in a single-pass mode, significantly reducing both time and space complexity. The core of our methodological innovation lies in an adaptive stochastic gradient descent algorithm tailored for dynamic objective functions, coupled with a novel online debiasing procedure. This allows us to maintain low-dimensional summary statistics while effectively controlling optimization errors introduced by the dynamically changing loss functions. We demonstrate, both methodologically and theoretically, that our method, termed the Approximated Debiased Lasso (ADL), not only mitigates the need for the bounded individual probability condition but also significantly improves numerical performance. Numerical experiments demonstrate that the proposed ADL method consistently exhibits robust performance across various covariance matrix structures.
Keywords: Confidence interval, LASSO, one-pass algorithm, stochastic gradient descent.
1 Introduction
Online statistical inference refers to methods and algorithms used for making inference about population parameters in a real-time environment, where data is collected sequentially rather than being static as in conventional settings. More specifically, these methods update parameter estimates and quantify their uncertainty along with sequentially accrued data information. As opposed to offline learning techniques that train over a static big dataset, online learning algorithms are expected to be computationally more efficient while achieving similar statistical performances to their offline counterparts. Most existing works in the field of online statistical inference focus on low-dimensional settings where the sample size ($n$) is much larger than the number of parameters ($p$). Their main idea is to save and update a $p$ by $p$ information matrix for quantifying the uncertainty in parameter estimation (Schifano et al.,, 2016; Luo and Song,, 2020; Luo et al., 2023b, ). Nevertheless, this idea is not directly extendable to high-dimensional settings where $p\gg n$ because the space complexity of saving such a huge matrix far exceeds that of saving the raw data. It is worth noting that most existing work in online high-dimensional settings either requires the availability of an entire raw dataset or the storage of summary statistics of large dimensions (Shi et al.,, 2020; Deshpande et al.,, 2023; Luo et al., 2023a, ).
In online inference, one of the main reasons to store the information matrix is that this quantity is pivotal in statistical inference and one can use such information directly to construct confidence interval for the target parameter(s). Moreover, online algorithms that leverage this second-order information can provide more accurate direction and step size towards the minimum, learning to faster convergence in some scenarios (Toulis and Airold,, 2017). However, as mentioned earlier, storing a $p\times p$ information matrix is generally not preferred in high-dimensional settings. As an alternative approach, the stochastic gradient descent (SGD) algorithm (Robbins and Monro,, 1951) fits the online setting well due to its computational and memory efficiency. In particular, SGD only requires one pass over the data, and its space complexity can reach $\mathcal{O}(p)$. The pioneering work that uses SGD for statistical inference can be traced back to Polyak and Juditsky, (1992) and it has also been applied to the low-dimensional models thereafter (Toulis and Airold,, 2017; Fang,, 2019). Recently, Chen et al., (2020) incorporate the SGD into high-dimensional inference task. Specifically, they combine the regularization annealed epoch dual averaging (RADAR) algorithm, one variant of SGD (Agarwal et al.,, 2012), with debiasing techniques (Zhang and Zhang,, 2014; van de Geer et al.,, 2014) in the linear models. However, their method requires access to the entire dataset in the debiasing step. Lately, Han et al., (2024) propose a new online debiased SGD algorithm in high-dimensional linear models that is a one-pass algorithm with space complexity $\mathcal{O}(p)$. While this algorithm has many desirable properties, it is not directly applicable to high-dimensional nonlinear models such as generalized linear models. In summary, to the best of our knowledge, existing online inference methods for high-dimensional generalized linear models are either not one-pass or require a space complexity of $\mathcal{O}(p^{2})$.
The goal of this paper is to develop an online statistical inference method for high-dimensional generalized linear models. In particular, the proposed method is able to sequentially update the regression coefficient estimates and their standard errors upon the arrival of a new data point, without retrieving historical raw data of space complexity $\mathcal{O}(np)$ or saving summary statistics of space complexity $\mathcal{O}(p^{2})$. It is worth mentioning that the generalized linear model is quite different from the linear model, thus the methods in Chen et al., (2020) and Han et al., (2024) are not easily extendable to generalized linear models. The major challenges come from two aspects. First, the objective function in the online inference framework dynamically changes as the algorithm proceeds forward. This is fundamentally different from the RADAR algorithm where the objective function is fixed. As a result, the change in the objective function introduces an additional error term which should be carefully controlled to establish asymptotic normality. Second, due to the nonlinear nature of the generalized linear models, it is nontrivial to develop a one-pass algorithm that utilizes summary statistics of low-space complexity. In both the first-order gradient and Hessian matrix of the loss function, parameter estimates and data are intertwined together, making a direct decomposition infeasible.
To address these challenges, we propose a new online debiasing procedure with two key components. First, we propose an adaptive RADAR algorithm that allows the underlying objective function change dynamically. More importantly, the optimization errors induced from the dynamically changed objective functions is still controllable, which provides the basis to further conduct valid statistical inference. Second, to avoid reaccessing and recomputing with historical raw data, we propose to linearize the nonlinear form of the first-order gradient via Taylor approximation. Such a linearization step is highly non-trivial. It is clear that second-order approximation can effectively reduce approximation errors, but it will involve Hessian matrix with space complexity of $\mathcal{O}(p^{2})$. Our key idea is to construct the summary statistics by combining the output from the proposed adaptive RADAR algorithm and the second-order information to reduce the dimensionality. In particular, these summary statistics require only $\mathcal{O}(p)$ space complexity but still retain the key information to conduct inference.
Our contributions can be summarized from both methodological and theoretical aspects. Methodologically, we propose an online inference method for high-dimensional inference, named the approximated debiased lasso (ADL). The proposed method offers two main computational benefits. First, it is designed to operate in a single-pass mode, resulting in low time complexity. Second, this method does not require storing the raw data or large summary statistics, and its space complexity is only of $\mathcal{O}(p)$, which is comparable to the complexity of storing a single data point. Thus, our method can handle ultra-high dimensional data effectively. In our real data analysis, we demonstrate that our method can handle datasets with millions of covariates on a standard personal computer. By contrast, alternative methods requiring access to the entire raw dataset or large summary statistics fail in similar scenarios due to memory constraints. A detailed comparison between our methods and relevant existing literature is presented in Table 1.
In our theoretical work, we have removed the bounded individual probability condition – a condition commonly imposed in the literature to establish the asymptotic theorem in high-dimensional settings. This condition typically applies constraints to the sample data. Instead, our approach shifts focus from statistical error to optimization error, especially errors arising from the SGD algorithm. Notably, the objective function for SGD is based on the loss function at the population level rather than the sample level, so that we can avoid making assumptions on the sample data. We believe this SGD approach offers a novel perspective to handle many other statistical problems that are of independent interest. In addition, removing such a technical condition is not only a theoretical advancement but also improves the numerical performance. As demonstrated in Cai et al., (2023), the bounded individual probability condition is easily violated when the variance of the covariate is large, deteriorating the performance of most conventional high-dimensional inference methods. In contrast, our method consistently performs well even with varying covariance matrix structures, demonstrating robustness under different scenarios.
Table 1: The time and space complexity of conducting inference for a single parameter in high-dimensional models.
Work Model One-pass? Space complexity
Chen et al., (2020) LM Yes $\mathcal{O}(np)$
Deshpande et al., (2023) LM No $\mathcal{O}(p^{2})$
Han et al., (2024) LM Yes $\mathcal{O}(p)$
Shi et al., (2020) GLM No $\mathcal{O}(np)$
Luo et al., 2023a GLM No $\mathcal{O}(p^{2})$
ADL (Proposed Method) GLM Yes $\mathcal{O}(p)$
Notes: LM refers to linear model; GLM refers to generalized linear model.
Notation: We use $\vee$ and $\wedge$ to denote the $max$ and $min$ operator, respectively. For a random variable $Z$, its sub-Gaussian norm is defined as $\lVert Z\lVert_{\psi_{2}}=\sup_{q\geq 1}q^{-1/2}(\mathbb{E}|Z|^{q})^{1/q}$; For a random vector $\bm{Z}$, its sub-Gaussian norm is defined as $\lVert\bm{Z}\lVert_{\psi_{2}}=\sup_{\lVert\bm{v}\lVert_{2}=1}\lVert\bm{v}^{T}\bm{Z}\lVert_{\psi_{2}}$. We say $\bm{Z}$ is sub-Gaussian if its sub-Gaussian $\lVert\bm{Z}\lVert_{\psi_{2}}$ norm is finite. We also call a constant $c$ the universal constant if $c$ does not depend on $n$ and $p$. For $m\in\mathbb{N}$, we write $\{1,\cdots,m\}$ as $[m]$. We use $\mathcal{O}_{p}(\cdot)$ and $o_{p}(\cdot)$ to represent the equivalent and higher-order relation between two random sequences in probability, respectively. For a deterministic sequence, we use $\mathcal{O}(\cdot)$ in place of $\mathcal{O}_{p}(\cdot)$. We use $\lesssim$ and $\gtrsim$ to denote the asymptotic inequality relations.
2 Methodology
Suppose that $n$ data points $\{(\bm{x}_{i},y_{i})\}_{i=1}^{n}$ arrive sequentially and each of them is an independent and identically distributed copy of $(\bm{X},Y)$, which are sampled from a generalized linear model:
$\mathbb{P}(Y\lvert\bm{X};\bm{\beta}^{*})\propto\exp\left\{\frac{Y(\bm{X}^{\intercal}\bm{\beta}^{*})-\Phi(\bm{X}^{\intercal}\bm{\beta}^{*})}{c(\sigma)}\right\},$ (1)
where $\Phi(\cdot)$ is a link function, $c(\sigma)$ is a fixed and known parameter, and $\bm{\beta}^{*}$ is an $s_{0}$-sparse parameter, that is $\lVert\bm{\beta}^{*}\lVert_{0}=s_{0}$. Our aim is to provide a confidence interval estimator of a single coefficient or low-dimensional regression coefficients in $\bm{\beta}^{*}$ along with sequentially collected data points.
In the offline setting, one of the most popular approaches is the debiased lasso method (Zhang and Zhang,, 2014; van de Geer et al.,, 2014). The debiased lasso method consists of two steps: obtain the lasso estimator and correct the bias in the lasso estimator. Under the generalized linear model, we denote the negative log-likelihood function as $F_{n}(\bm{\beta})=({1}/{n})\sum_{i=1}^{n}\left\{-y_{i}(\bm{x}^{\intercal}_{i}\bm{\beta})+\Phi(\bm{x}^{\intercal}_{i}\bm{\beta})\right\}$, and specify its $\ell_{1}$-penalized form as follows:
$F_{n}(\bm{\beta};\lambda_{n})=F_{n}(\bm{\beta})+\lambda_{n}\lVert\bm{\beta}\lVert_{1}.$ (2)
Then the lasso estimator is defined as $\widetilde{\bm{\beta}}^{(n)}=\arg\min_{\bm{\beta}\in\mathbb{R}^{p}}F_{n}(\bm{\beta};\lambda_{n}).$ Since $\widetilde{\bm{\beta}}^{(n)}$ does not have a tractable limiting distribution (Zhao and Yu,, 2006), the debiased step was considered. Specifically, for some $j\in[p]$, the debiased estimator is calculated via
$\widetilde{\bm{\beta}}^{(n)}_{j,de}=\widetilde{\bm{\beta}}_{j}^{(n)}-\{\widetilde{\bm{\gamma}}^{(n)}_{\cdot,j}\}^{\intercal}f_{n}(\widetilde{\bm{\beta}}^{(n)}),$ (3)
where $f_{n}(\widetilde{\bm{\beta}}^{(n)})=\nabla F_{n}(\widetilde{\bm{\beta}}^{(n)})$ is the gradient of the negative log-likelihood function, and $\widetilde{\bm{\gamma}}^{(n)}_{\cdot,j}$, known as the the nodewise lasso estimator, serves as an estimator of the $j$-th column in the inverse of Hessian matrix $\nabla^{2}F_{n}(\widetilde{\bm{\beta}}^{(n)})$. Zhang and Zhang, (2014) and van de Geer et al., (2014) show that such a debiased estimator converges to the normal distribution under the linear model and generalized linear model respectively.
It is straightforward to check whether (3) could fit into the online learning framework. Apparently, the calculations of $\widetilde{\bm{\beta}}^{(n)}$, $\widetilde{\bm{\gamma}}^{(n)}_{\cdot,j}$, and $f_{n}(\widetilde{\bm{\beta}}^{(n)})$ all require re-accessing and recomputing with the full dataset. For this reason, we consider developing real online methods and algorithms that update the debiased lasso estimator without revisiting historical data. Based on existing literature in online lasso estimation, online alternative for estimating $\widetilde{\bm{\beta}}^{(n)}$ is readily available (Agarwal et al.,, 2012). Nonetheless, as indicated by (3), there are two remaining components we need to deal with to conduct a one-step estimation that fits into an online statistical inference framework.
In the following, we develop an online method to conduct statistical inference. We first introduce the RADAR algorithm used to conduct online lasso estimation in Section 2.1. Specifically, we develop an adaptive RADAR algorithm to find online alternatives to replace $\widetilde{\bm{\gamma}}^{(n)}_{\cdot,j}$ in Section 2.2. We then propose an approximated debiasing technique to replace $f_{n}(\widetilde{\bm{\beta}}^{(n)})$. The resulting debiased estimator, as well as its standard error, are presented in Sections 2.32.4. For illustration, we present the entire procedure in Figure 1. The online lasso estimation algorithm is marked in blue. Our proposed method, including the adaptive RADAR and approximated debiasing technique, is highlighted in orange. We wish to emphasize that our procedure is one-pass and only requires $\mathcal{O}(p)$ space complexity.
Figure 1: The whole procedure in conducting online statistical inference.
2.1 Online lasso with RADAR
In this subsection, we introduce an online one-pass approach to finding the lasso estimator without retrieving or recomputing with historical data. A natural choice that serves this purpose is the SGD (Robbins and Monro,, 1951). Although SGD can obtain an estimator of $\bm{\beta}^{*}$, its convergence rate is unsatisfactory in high-dimensional settings. For example, Han et al., (2024) show that interval estimation using the vanilla SGD may not have desirable performance such as attaining nominal coverage probability. To address this issue, we adopt the regularization annealed epoch dual averaging (RADAR) algorithm (Agarwal et al.,, 2012), a variant of the vanilla SGD.
Different from the vanilla SGD, the RADAR algorithm incorporates the multi-epoch technique. The multi-epoch technique splits a sequence of $n$ data points into $K$ epochs according to a sequence of epoch length schedule $\{T_{k}\}_{k\in[K]}$, where $k$-th epoch contains data of sample size $T_{k}$ and $\sum_{k=1}^{K}T_{k}=n$. Let $n_{k}=\sum_{i=1}^{k}T_{i}$ denote the cumulative sample size up to the $k$-th epoch. In the $k$-th epoch, the RADAR algorithm runs $T_{k}$ rounds of the proximal SGD with dual averaging (Xiao,, 2009) and outputs the estimate $\widehat{\bm{\beta}}^{(n_{k})}$ at the end of the epoch. The output $\widehat{\bm{\beta}}^{(n_{k})}$ is then used as an initial value in the $(k+1)$-th epoch. The multi-epoch method accelerates the convergence rate by dropping the information received in the previous epochs. Since the RADAR algorithm outputs the estimate only at the end of each epoch, $\widehat{\bm{\beta}}^{(i)}$ is well-defined only for $i=n_{k},k\in[K]$. For convenience, we define the estimates within the $(k+1)$-th epoch as
$\widehat{{\bm{\beta}}}^{(i)}=\widehat{{\bm{\beta}}}^{(n_{k})},\ \text{if}\ n_{k}+1\leq i<n_{k+1}.$ (4)
In other words, the estimates within each epoch remain unchanged until the last observation in this epoch arrives. Based on (4), we obtain a sequence of the online lasso estimators, $\{\widehat{{\bm{\beta}}}^{(i)}\}_{i\geq n_{1}}$ by the RADAR algorithm.
It is worth noting that $\widehat{{\bm{\beta}}}^{(n)}$, the final output from the RADAR algorithm, is no longer the minimizer of $F_{n}(\bm{\beta};\lambda_{n})$ defined in (2). In fact, without accessing the entire dataset, it is hard to achieve the optimality as the offline estimator. Recall $\widetilde{\bm{\beta}}^{(n)}$ is the minimizer of $F_{n}(\bm{\beta};\lambda_{n})$. Instead of the statistical error $\lVert\widetilde{\bm{\beta}}^{(n)}-\bm{\beta}^{*}\lVert_{1}$, we need to consider the optimization error from RADAR, that is $\lVert\widehat{{\bm{\beta}}}^{(n)}-\bm{\beta}^{*}\lVert_{1}$. Under some regularity conditions, we can conclude that the optimization error $\lVert\widehat{{\bm{\beta}}}^{(n)}-\bm{\beta}^{*}\lVert_{1}$ achieves the same rate as the statistical error $\lVert\widetilde{\bm{\beta}}^{(n)}-\bm{\beta}^{*}\lVert_{1}$. A detailed description of the RADAR algorithm is included in the supplementary material.
2.2 Online nodewise lasso with adaptive RADAR
Given the online lasso estimator, we still need to find online alternatives to $\widetilde{\bm{\gamma}}^{(n)}_{\cdot,j}$ and $f_{n}(\widetilde{\bm{\beta}}^{(n)})$ defined in (3), for the one-step estimation. In an offline setting, the former is computed via the nodewise lasso algorithm, that is,
$\operatorname*{arg\,min\;}_{\bm{r}\in\mathbb{R}^{{p}},\bm{r}_{j}=-1}\frac{1}{2n}\sum_{i=1}^{n}\left[\ddot{\Phi}(\bm{x}_{i}^{\intercal}\widetilde{\bm{\beta}}^{(n)})\{(\bm{x}_{i})^{\intercal}\bm{r}\}^{2}\right]+\lambda_{n}^{\prime}\lVert\bm{r}\lVert_{1}.$ (5)
To solve (5) in an online fashion, one may again adopt the SGD approach, such as the RADAR algorithm. However, according to (5), the stochastic gradient with respect to $\bm{r}_{-j}$ contains the term $(\bm{x}_{i})_{-j}\{(\bm{x}_{i})^{\intercal}\bm{r}\}\ddot{\Phi}(\bm{x}_{i}^{\intercal}\widetilde{\bm{\beta}}^{(n)}),\ i\in[n]$, where $\widetilde{\bm{\beta}}^{(n)}$ is not available when $i<n$. Thus, the RADAR algorithm is not directly applicable to solving the nodewise lasso problem.
To address this issue, we modify the RADAR algorithm by replacing $\widetilde{\bm{\beta}}^{(n)}$ with the latest online lasso estimator of $\bm{\beta}^{*}$. Since the resulting gradient is adaptive to the sequentially changed lasso estimator of $\bm{\beta}^{*}$, we call this algorithm the adaptive RADAR algorithm. Note that the initial estimator of $\bm{\beta}^{*}$ is only available after processing the first $n_{1}$ data. Thus, we start the online nodewise lasso after the first $n_{1}$ data points. In particular, we consider the epoch length $\{T_{k}^{\prime}\}_{k\in[K^{\prime}]}$ and define the cumulative length or sample size as $n_{k}^{\prime}=n_{1}+\sum_{j=1}^{k}T_{k}^{\prime}$ for $k\in[K^{\prime}]$. Our newly proposed approach computes the stochastic gradient of $\bm{r}_{-j}$ via
$(\bm{x}_{i})_{-j}\{(\bm{x}_{i})^{\intercal}\bm{r}\}\ddot{\Phi}(\bm{x}_{i}^{\intercal}\widehat{\bm{\beta}}^{(n_{k-1}^{\prime})})+\lambda_{k}^{\prime}\textit{sign}(\bm{r}_{-j}),\ \ \ n_{k-1}^{\prime}<i\leq n_{k}^{\prime},$ (6)
in the $k$-th epoch where $\widehat{\bm{\beta}}^{(n_{k-1}^{\prime})}$ is defined in (4) and sign is the sign function. Then, we adopt (6) as the stochastic gradient and follow the rest of the procedure in the RADAR algorithm. As a result, we obtain the output at the end of $k$-th epoch, denoted by ${\widehat{\bm{\gamma}}_{\cdot,j}}^{(n_{k}^{\prime})}$ for $k\in[K^{\prime}]$. Similar to (4), we also define
${\widehat{\bm{\gamma}}_{\cdot,j}}^{(i)}={\widehat{\bm{\gamma}}_{\cdot,j}}^{(n_{k}^{\prime})},\ \text{if}\ n_{k}^{\prime}\leq i<n_{k+1}^{\prime},$ (7)
for $k\in[K^{\prime}-1]$. Then, we treat ${\widehat{\bm{\gamma}}_{\cdot,j}}^{(i)},\ i\geq n_{1}^{\prime}$ as an online nodewise lasso estimator.
The adaptive RADAR is more general than the original RADAR algorithm. Specifically, let $\mathcal{F}_{k}^{\prime}=\sigma\left(\{(\bm{x}_{i},y_{i})\}_{i=1}^{n_{k}^{\prime}}\right)$, that is the $\sigma$-field generated from $\{(\bm{x}_{i},y_{i})\}_{i=1}^{n_{k}^{\prime}}$, and $\bar{G}_{k}(\bm{r})=\mathbb{E}\left(\ddot{\Phi}(\bm{X}^{\intercal}\widehat{\bm{\beta}}^{(n_{k-1}^{\prime})})(\bm{X}^{\intercal}\bm{r})^{2}\big{\lvert}\mathcal{F}_{k-1}^{\prime}\right)$. In the $k$-th epoch, the stochastic gradient defined in (6) comes from the objective function
$\bar{G}_{k}(\bm{r})+\lambda_{k}^{\prime}\lVert\bm{r}\lVert_{1},$ (8)
which contains two parts: the loss function $\bar{G}_{k}(\bm{r})$ and the regularization term $\lambda_{k}^{\prime}\lVert\bm{r}\lVert_{1}$. We wish to highlight that both terms vary across different epochs. In contrast, the objective function in the original RADAR algorithm (Agarwal et al.,, 2012) is
$G(\bm{r})+\lambda_{k}^{\prime}\lVert\bm{r}\lVert_{1}.$ (9)
for some fixed loss function $G$. Compared to (9), an innovative aspect of our proposed objective function (8) is that, the loss function can vary as the optimization algorithm proceeds, though it brings new challenge in the theoretical justification. In particular, changing the loss function also changes the target value of the optimization problem, which induces additional errors. With a suitable design of the hyperparameters, we show that this approach will not have a larger bias and the overall error can still be controlled at a convergence rate of $\mathcal{O}(1/\sqrt{n})$ under some regularity conditions, see Theorem 3.2.
Remark 1.
We would like to comment on the difference between SGD and gradient descent in an online setting. Briefly speaking, they optimize over different objective functions. The gradient descent aims to solve an optimization problem where the loss function is calculated based on random samples, for example, $F_{n}(\bm{\beta};\lambda_{n})$ in (2). In contrast, the (online) SGD focuses on the loss function at the population level, such as $\mathbb{E}\{F_{n}(\bm{\beta};\lambda_{n})\}$. This main difference calls for the development of different techniques in the theoretical justification. For example, the restricted strong convexity condition of the loss function is one of the key conditions to establish the oracle inequality in the lasso regression. However, in the (online) SGD, we impose the strong convexity of the population loss function rather than the empirical loss function. As a result, our theoretical justification does not involve those conditions for handling the sample loss function, such as the bounded individual probability condition (van de Geer et al.,, 2014): $|\bm{x}_{i}^{\intercal}\bm{\beta}^{*}|<M,i\in[n]$ for a universal constant $M.$
2.3 Approximated debiasing techniques
In generalized linear models, computing $f_{n}(\widehat{\bm{\beta}}^{(n)})$ is infeasible in an online setting. Specifically, $f_{n}(\widehat{\bm{\beta}}^{(n)})=\nabla F_{n}(\widehat{\bm{\beta}}^{(n)})$, taking the following form
$f_{n}(\widehat{\bm{\beta}}^{(n)})=\frac{1}{n}\sum_{i=1}^{n}\bm{x}_{i}\left\{-y_{i}+\dot{\Phi}(\bm{x}^{\intercal}_{i}\widehat{\bm{\beta}}^{(n)})\right\}.$ (10)
Due to the non-linearity of the function $\dot{\Phi}(\cdot)$, the computation of $\sum_{i=1}^{n}\bm{x}_{i}\dot{\Phi}(\bm{x}^{\intercal}_{i}\widehat{\bm{\beta}}^{(n)})$ in (10) is impractical if we cannot retrieve the historical data $\{\bm{x}_{i}\}_{i<n}$ after obtaining $\widehat{\bm{\beta}}^{(n)}$. For instance, in the logistics regression $\Phi(t)=\{1/(1+\exp(-t))\}$, there is no such a summary statistics to obtain $\sum_{i=1}^{n}\bm{x}_{i}\dot{\Phi}(\bm{x}^{\intercal}_{i}\widehat{\bm{\beta}}^{(n)})$.
Instead of computing the exact value of $\sum_{i=1}^{n}\bm{x}_{i}\dot{\Phi}(\bm{x}^{\intercal}_{i}\widehat{\bm{\beta}}^{(n)})$, we will approximate it. A straightforward solution is the plug-in approximation, namely $\sum_{i=1}^{n}\bm{x}_{i}\dot{\Phi}(\bm{x}^{\intercal}_{i}\widehat{\bm{\beta}}^{(i)})$. However, complications arise from the approximation error from the plug-in method will involve one additional term whose order is $\mathcal{O}(\sqrt{n}).$ Consequently, we cannot derive the asymptotic normality. To circumvent the problem, we employ the approximated debiasing techniques via Taylor approximation. Specifically,
$\dot{\Phi}(\bm{x}^{\intercal}_{i}\widehat{\bm{\beta}}^{(n)})=\dot{\Phi}(\bm{x}^{\intercal}_{i}\widehat{\bm{\beta}}^{(i)})+\ddot{\Phi}(\bm{x}^{\intercal}_{i}\widehat{\bm{\beta}}^{(i)})\bm{x}^{\intercal}_{i}(\widehat{\bm{\beta}}^{(n)}-\widehat{\bm{\beta}}^{(i)})+\mathcal{O}(\lVert\widehat{\bm{\beta}}^{(n)}-\widehat{\bm{\beta}}^{(i)}\lVert_{2}^{2}),\ i\in[n].$ (11)
The key idea of our method is to use the second-order approximation rather than the plug-in approximation, namely,
$\sum_{i=1}^{n}\bm{x}_{i}\dot{\Phi}(\bm{x}^{\intercal}_{i}\widehat{\bm{\beta}}^{(n)})\approx\sum_{i=1}^{n}\left\{\bm{x}_{i}\dot{\Phi}(\bm{x}^{\intercal}_{i}\widehat{\bm{\beta}}^{(i)})+\bm{x}_{i}\ddot{\Phi}(\bm{x}^{\intercal}_{i}\widehat{\bm{\beta}}^{(i)})\bm{x}^{\intercal}_{i}(\widehat{\bm{\beta}}^{(n)}-\widehat{\bm{\beta}}^{(i)})\right\}.$
Note that the Hessian matrices $\bm{x}_{i}\ddot{\Phi}(\bm{x}^{\intercal}_{i}\widehat{\bm{\beta}}^{(i)})\bm{x}^{\intercal}_{i},\ i\in[n]$ are $p\times p$ matrices, which requires a large amount of memory capacity. To ensure the superiority of our method in terms of the space complexity, we construct the following summary statistics. Letting $l=\min\{L:n_{L}\geq n_{1}^{\prime}\}$, we compute
$\bm{a}_{1}^{(k)}=\sum_{i=n_{l}+1}^{k}\bm{x}_{i}\left\{-y_{i}+\dot{\Phi}(\bm{x}^{\intercal}_{i}\widehat{\bm{\beta}}^{(i)})\right\},\ \bm{a}_{2}^{(k)}=\sum_{i=n_{l}+1}^{k}\bm{x}_{i}^{\intercal}{\widehat{\bm{\gamma}}_{\cdot,j}}^{(i)}\ddot{\Phi}(\bm{x}^{\intercal}_{i}\widehat{\bm{\beta}}^{(i)})\bm{x}_{i},$
$a_{3}^{(k)}=\sum_{i=n_{l}+1}^{k}\bm{x}_{i}^{\intercal}{\widehat{\bm{\gamma}}_{\cdot,j}}^{(i)}\ddot{\Phi}(\bm{x}^{\intercal}_{i}\widehat{\bm{\beta}}^{(i)})\bm{x}^{\intercal}_{i}\widehat{\bm{\beta}}^{(i)},\ {a}_{4}^{(k)}=\sum_{i=n_{l}+1}^{k}\bm{x}_{i}^{\intercal}{\widehat{\bm{\gamma}}_{\cdot,j}}^{(i)}(\bm{x}_{i})_{j}\left\{\ddot{\Phi}(\bm{x}_{i}^{\intercal}\widehat{\bm{\beta}}^{(i)})\right\},$ (12)
where $\bm{a}_{1}^{(k)}$ and $\bm{a}_{2}^{(k)}$ are $p$-dimensional vectors while the other elements $a_{3}^{(k)}$ and $a_{4}^{(k)}$ are scalars. Note that the above summary statistics are computed from $i=n_{l}+1$ rather than $i=1$. The main purpose is to guarantee the existence of ${\widehat{\bm{\gamma}}_{\cdot,j}}^{(i)}$ and $\widehat{\bm{\beta}}^{(i)}$.
Based on the statistics defined in (12), we define the approximated debiased lasso (ADL) estimator as
${\widehat{\bm{\beta}}}^{(k)}_{j,\text{de}}:=\widehat{{\bm{\beta}}}^{(k)}_{j}-\frac{\{\bm{a}^{(k)}_{1}\}^{\intercal}{\widehat{\bm{\gamma}}_{\cdot,j}}^{(k)}+\{\bm{a}^{(k)}_{2}\}^{\intercal}\widehat{{\bm{\beta}}}^{(k)}-a_{3}^{(k)}}{{a}_{4}^{(k)}},\ k>n_{l}.$ (13)
One key advantage of (13) is that we can update these summary statistics recursively instead of storing the raw data in memory. For example, when the $k$-th data arrives, we compute
$\bm{a}_{1}^{(k)}=\bm{a}_{1}^{(k-1)}+\bm{x}_{k}\left\{-y_{k}+\dot{\Phi}(\bm{x}^{\intercal}_{k}\widehat{\bm{\beta}}^{(k)})\right\}.$ (14)
Then, following the update as in (14), the calculation of ${\widehat{\bm{\beta}}}^{(k)}_{j,\text{de}}$ in (13) is one-pass with $\mathcal{O}(p)$ space complexity. For ease of illustration, a diagram that depicts the alignment of the three key steps in online estimation and inference procedure is presented in Figure 2.
Figure 2: The alignment of three key steps in constructing the ADL estimator.
2.4 Variance and confidence interval
To construct confidence interval, we still need to find the variance of ${\widehat{\bm{\beta}}}^{(k)}_{j,\text{de}}$. To this end, we design another scalar-valued summary statistic $a_{5}^{(k)}=\sum_{i=n_{l}+1}^{k}(\bm{x}_{i}^{\intercal}{\widehat{\bm{\gamma}}_{\cdot,j}}^{(i)})^{2}\left\{\dot{\Phi}(\bm{x}^{\intercal}_{i}\widehat{\bm{\beta}}^{(i)})-y_{i}\right\}^{2}$. Then the estimated standard error of the ADL estimator is computed as
$\widehat{\nu}^{(k)}_{j}=\sqrt{a_{5}^{(k)}}/|a_{4}^{(k)}|.$ (15)
Consequently, based on (13) and (15), the $(1-\alpha)$-confidence interval for $\bm{\beta}_{j}^{*}$ with $10<\alpha<$ 1 at time $k$ is given by:
$\left({\widehat{\bm{\beta}}}^{(k)}_{j,\text{de}}-z_{\alpha/2}\widehat{\nu}^{(k)}_{j},{\widehat{\bm{\beta}}}^{(k)}_{j,\text{de}}+z_{\alpha/2}\widehat{\nu}^{(k)}_{j}\right),\ k>n_{l},$ (16)
where $z_{\alpha/2}$ is the $\{1-(\alpha/2)\}$-quantile of the standard normal distribution. We present the pseudocode of the approximated debiasing procedure in Algorithm 1. Algorithm 1 processes each data point only once and does not store any original raw data. In addition, the required summary statistics are at most $p$-dimensional. Therefore, Algorithm 1 is one-pass and only requests $\mathcal{O}(p)$ space complexity, as stated in Table 1.
Input : Streaming data $\{(\bm{x}_{i},y_{i})\}_{i=1}^{n}$, epoch schedule $\{T_{k}\}_{k\in[K]}$ and $\{T_{k}^{\prime}\}_{k\in[K^{\prime}]}$, regularization parameters $\{\lambda_{k}\}_{k\in[K]}$ and $\{\lambda_{k}^{\prime}\}_{k\in[K^{\prime}]}$, significance level $\alpha$, and index of variable of interest $j$.
1 Initialize $\widehat{\bm{\beta}}^{(0)},{\widehat{\bm{\gamma}}_{\cdot,j}}^{(0)},\bm{a}_{1}^{(0)},\bm{a}_{2}^{(0)},a_{3}^{(0)},a_{4}^{(0)},a_{5}^{(0)}$;
2 for $i=1,\ldots,n$ do
3      Receive the data point $(\bm{x}_{i},y_{i})$;
4       Conduct the RADAR with the hyper-parameters $\{T_{k},\lambda_{k}\}_{k\in[K]}$ ;
5       Conduct the adaptive RADAR with the hyper-parameters $\{T_{k}^{\prime},\lambda_{k}^{\prime}\}_{k\in[K^{\prime}]}$ ;
6       Obtain the online estimator $\widehat{\bm{\beta}}^{(i)}$ and ${\widehat{\bm{\gamma}}_{\cdot,j}}^{(i)}$ via (4) and (7) respectively;
7       Update $\bm{a}_{1}^{(i)},\bm{a}_{2}^{(i)},a_{3}^{(i)},a_{4}^{(i)},a_{5}^{(i)}$ recursively as (14);
8       Compute ${\widehat{\bm{\beta}}}^{(i)}_{j,\text{de}}$ and $\widehat{\nu}^{(i)}_{j}$ via (13) and (15) respectively;
9       Output the $(1-\alpha)$-confidence interval
10                 $\left({\widehat{\bm{\beta}}}^{(i)}_{j,\text{de}}-z_{\alpha/2}\widehat{\nu}^{(i)}_{j},{\widehat{\bm{\beta}}}^{(i)}_{j,\text{de}}+z_{\alpha/2}\widehat{\nu}^{(i)}_{j}\right)$ ;
11       Update $\left(\widehat{\bm{\beta}}^{(i)},{\widehat{\bm{\gamma}}_{\cdot,j}}^{(i)},\bm{a}_{1}^{(i)},\bm{a}_{2}^{(i)},a_{3}^{(i)},a_{4}^{(i)},a_{5}^{(i)}\right)$ and clear the raw data in memory;
12 end for
Algorithm 1 Pseudocode of the proposed ADL algorithm.
3 Theoretical Properties
3.1 Oracle inequalities
To show the consistency of the RADAR output $\widehat{\bm{\beta}}^{(n)}$, we need the following assumptions.
Assumption 1.
The predictor vector $\bm{X}$ is a zero-mean and sub-Gaussian random vector, whose sub-Gaussian norm is $\kappa_{1}$ and the covariance matrix $\bm{\Sigma}$ satisfies: $0<M_{1}^{-1}\leq\Lambda_{\min}(\bm{\Sigma})\leq\Lambda_{\max}(\bm{\Sigma})\leq M_{1}<\infty$ for a universal constant $M_{1}$.
Assumption 2.
The response $Y$ is a sub-Gaussian random variable with the sub-Gaussian norm $\kappa_{2}$. In addition, the second derivative of the link function is Lipschitz and uniformly bounded by a constant $M_{\Phi}$, that is, $\max_{t\in\mathbb{R}}\lvert\ddot{\Phi}(t)\lvert\leq M_{\Phi}$.
Assumption 3.
Assume $\lVert\bm{\beta}^{*}\lVert_{0}=s_{0}$ and $\lVert\bm{\beta}^{*}\lVert_{2}=m$ for a universal constant $m$.
Assumption 1 considers the sub-Gaussian random design, which is commonly adopted in establishing the oracle inequality for the high-dimensional model in the literature. Assumption 2 is a mild regularity condition on the generalized linear model (1). It is satisfied by a broad class of models, such as linear model with sub-Gaussian error, logistic model, and probit model. Assumption 3 is a basic assumption on the parameter space in the sparse regression settings.
Let $d_{0}=\lVert\widehat{\bm{\beta}}^{(0)}-\bm{\beta}^{*}\lVert_{1}$ denote the initial error of the algorithm. Without loss of generality, we assume that $d_{0}\gtrsim 1.$ Then, the following theorem establishes the rate of convergence of the output from the RADAR algorithm.
Theorem 3.1.
Suppose that Assumptions 1, 2 and 3 hold. With the choice of hyper-parameters in (20), the following events
$\lVert\widehat{\bm{\beta}}^{(i)}-\bm{\beta}^{*}\lVert_{1}\leq C_{1}s_{0}d_{0}\sqrt{\frac{(\log p)^{3}}{i}},\ \ \lVert\widehat{\bm{\beta}}^{(i)}-\bm{\beta}^{*}\lVert_{2}\leq C_{1}d_{0}\sqrt{\frac{s_{0}(\log p)^{3}}{i}}$ (17)
for $i\geq n_{1}$ hold uniformly for a universal constant $C_{1}$ with probability at least $1-7(\log p)^{-6}.$
Compared to the offline approach (Negahban et al.,, 2012), our error bound has an additional term: the initial error $d_{0}$. This is because our online algorithm is one-pass and iterates through its input data exactly once. This means that the number of iterations would equal to the size of the input dataset which is finite in practice. With this regard, the initial error is not negligible and should be taken into consideration.
Remark 2.
If we revise the sub-Gaussian design in Assumption 1 to the uniformly bounded design, that is, $\lVert\bm{X}\lVert_{\infty}$ is bounded by some universal constant, the convergence rate in terms of $\ell_{1}$-norm will be $\mathcal{O}_{\mathbb{P}}(s_{0}d_{0}\sqrt{(\log p)/i})$ for $i\geq n_{1}$. But the uniformly bounded design contradicts our simulation settings where $\bm{X}$ follows the normal distribution. Hence, the theoretical result under the uniformly bounded design is not presented.
Next, we consider the estimation error of the output from the adaptive RADAR. Recall that the inverse of Hessian matrix is ${\bm{\Theta}}(\bm{\beta})=[\mathbb{E}\{\bm{X}^{\intercal}\ddot{\Phi}(\bm{X}^{\intercal}\bm{\beta})\bm{X}\}]^{-1}$ and its normalized version is $\overline{\bm{\Theta}}(\bm{\beta})=[\text{diag}\{{\bm{\Theta}}(\bm{\beta})\}]^{-1}{\bm{\Theta}}(\bm{\beta}).$ Two additional assumptions on the normalized inverse Hessian matrix are imposed.
Assumption 4.
There exists a universal constant $M_{2}$ such that $0<M_{2}^{-1}\leq\Lambda_{\min}(\overline{\bm{\Theta}}(\bm{\beta}^{*}))\leq\Lambda_{\max}(\overline{\bm{\Theta}}(\bm{\beta}^{*}))\leq M_{2}<\infty$. The columns of $\overline{\bm{\Theta}}(\bm{\beta}^{*})$ is $s_{1}$-sparse: $\max_{j\in[p]}\lVert\{\overline{\bm{\Theta}}(\bm{\beta}^{*})\}_{\cdot,j}\lVert_{0}$  $=s_{1}.$
Assumption 5.
There is a universal constant $L$ such that $\lVert\overline{\bm{\Theta}}(\bm{b}_{1})-\overline{\bm{\Theta}}(\bm{b}_{2})\lVert_{1}\leq L\lVert\bm{b}_{1}-\bm{b}_{2}\lVert_{1}$, for any $\bm{b}_{1},\bm{b}_{2}\in\{\bm{\beta}:\lVert\bm{\beta}-\bm{\beta}^{*}\lVert_{1}\leq 2d_{0}.\}$
Assumption 4 ensures the $\ell_{0}$-sparsity of $\bm{\gamma}_{\cdot,j}^{*}$ and the boundedness of $\lVert\bm{\gamma}_{\cdot,j}^{*}\lVert_{2}$ as in Assumption 3. This assumption is commonly adopted in establishing the oracle inequality of nodewise lasso, such as van de Geer et al., (2014), Belloni et al., (2016) and Ning and Liu, (2017). Assumption 5 requires the normalized inverse Hessian matrix to be a local Lipschitz operator, which is a new requirement compared with those in the literature of the nodewise lasso. Nonetheless, this assumption is similar to Assumption (D4) in van de Geer et al., (2014) to some extent. The former could be regarded as the population-level version of the latter. Note that Assumption 5 holds automatically under the high-dimensional linear model since $\overline{\bm{\Theta}}(\bm{\beta})$ degenerates to a constant matrix.
Similarly, let $d_{1}=\lVert{\widehat{\bm{\gamma}}_{\cdot,j}}^{(0)}-\bm{\gamma}_{\cdot,j}^{*}\lVert_{1}$ be the initial error. We provide the convergence rate of the output from the adaptive RADAR algorithm in the next theorem.
Theorem 3.2.
Suppose that Assumptions 15 holds. With the choice of hyper-parameters in (20) and (22), given any fixed $j\in[p]$, the following events
$\lVert{\widehat{\bm{\gamma}}_{\cdot,j}}^{(i)}-\bm{\gamma}_{\cdot,j}^{*}\lVert_{1}\leq C_{2}\left\{s_{1}d_{1}+(s_{1}+s_{0})d_{0}\right\}\sqrt{\frac{(\log p)^{3}}{i-n_{1}}}$ (18)
for $i\geq n_{1}^{\prime}$ hold uniformly for a universal constant $C_{2}$ with probability at least $1-14(\log p)^{-6}$.
In contrast to Theorem 3.1, the error bound in (18) not only involves the initial error $d_{1}$ and the sparsity $s_{1}$ of the inverse Hessian matrix, but also the error from the online lasso algorithm via $d_{0}$ and $s_{0}$. This phenomenon shows the difference between the RADAR algorithm and the adaptive RADAR algorithm. As shown in (8), the loss function in the adaptive RADAR algorithm is dynamically changing because of the most recent plug-in estimates from the online lasso procedure. In other words, the target values are instantly changing. Therefore, additional error will be introduced from the drifted target values. With careful design of the hyperparameters as in (22), we are able to control this additional error. Consequently, the overall error will converge at the rate of $\mathcal{O}(\sqrt{n})$.
3.2 Asymptotic normality
By Theorems 3.1 and 3.2, we can establish the asymptotic normality of ${\widehat{\bm{\beta}}}^{(k)}_{j,\text{de}}$ defined in (13).
Theorem 3.3.
Suppose that Assumptions 15 hold and
$(s_{0}s_{1}d_{0}d_{1}+s_{0}^{2}d_{0}^{2}+s_{0}s_{1}d_{0}^{2})(\log p)^{4}\log_{2}(n)=o(\sqrt{n}).$ (19)
With the choice of hyper-parameters in (20) and (22), given any fixed $j\in[p],$
$\frac{1}{\widehat{\nu}^{(n)}_{j}}({\widehat{\bm{\beta}}}^{(n)}_{j,\text{de}}-{{\bm{\beta}}}^{*}_{j})\to\mathcal{N}(0,1)\ \text{ in distribution}\ \ \text{ as }n\to\infty.$
Compared to existing works in high-dimensional inferences, our theoretical results are different in two aspects. On one hand, the conditions in Theorem 3.3 are made at the population level. In contrast, most of the existing inference methods (van de Geer et al.,, 2014; Ning and Liu,, 2017; Ma et al.,, 2021; Shi et al.,, 2020; Luo et al., 2023a, ) rely on the bounded individual probability condition. For example, in logistic regression, the bounded individual probability condition takes the form that $\mathbb{P}(y_{i}=1|\bm{x}_{i})\in(c,1-c)$ for $i\in[n]$ and some positive constant $c$. However, such a condition can be easily violated if $\bm{x}_{i}$ follows a normal distribution with a slightly large variance. To the best of our knowledge, except for Guo et al., (2021) and Cai et al., (2023), most existing methods require the bounded individual probability condition. Without the bounded individual probability condition, our method is shown to provide valid inference under the Gaussian design. The removal of such a condition is not only a technical advancement, but also a methodological improvement as our method is able to handle more general cases.
On the other hand, we require slightly stronger conditions in terms of the sparsity levels $s_{0}$ and $s_{1}$, sample size $n$, and the dimension of features $p$. This is the trade-off between online and offline algorithms. Specifically, the online estimator may not be the optimiser of some loss functions. Hence, some optimality conditions, such as the Karush–Kuhn–Tucker (KKT) condition, cannot be used. Therefore, we need to impose slightly stronger conditions compared to its offline counterpart so as to establish the asymptotic properties; see Theorem 4.11 in Deshpande et al., (2023) and Theorem 1 in Han et al., (2024).
We end this section by providing the validity of the confidence interval for $\bm{\beta}^{*}_{j}$ given in Algorithm 1 in the next corollary.
Corollary 3.4.
Under the conditions in Theorem 3.3, for any fixed $j\in[p]$ and $\alpha\in(0,1)$,
$\mathbb{P}\left\{\bm{\beta}^{*}_{j}\in\big{(}{\widehat{\bm{\beta}}}^{(n)}_{j,\text{de}}-z_{\alpha/2}\widehat{\nu}^{(n)}_{j},{\widehat{\bm{\beta}}}^{(n)}_{j,\text{de}}+z_{\alpha/2}\widehat{\nu}^{(n)}_{j}\big{)}\right\}\to 1-\alpha\ \text{ as }n\to\infty.$
4 Numerical studies
In this section, we conduct simulation studies to examine the finite-sample performance of the ADL in high-dimensional logistic regression. Besides that, we provide a real data example to show the ability of our method in dealing with ultra-high-dimensional online data. All experiments are performed with an 11th Gen Intel Core i7-11370H processor and 16GB memory storage.
4.1 Simulated data
We randomly generate a sequence of $n$ i.i.d. copies $\{(\bm{x}_{i},y_{i})\}_{i=1}^{n}$ realized from the logistic regression model $\mathbb{P}(y=1\lvert\bm{x};\bm{\beta}^{*})=1-\mathbb{P}(y=0\lvert\bm{x};\bm{\beta}^{*})={\exp(\bm{x}^{\intercal}\bm{\beta}^{*})}/\{1+\exp(\bm{x}^{\intercal}\bm{\beta}^{*})\}$, where $\bm{x}\sim\mathcal{N}(0,\bm{\Sigma})$. Recall that the cardinality of the support set of $\bm{\beta}^{*}$ is $s_{0}$. We divide the support set into two halves $\{D_{1},D_{2}\}$, with true coefficients $\bm{\beta}^{*}_{D_{1}}=1$ and $\bm{\beta}^{*}_{D_{2}}=-1$ respectively. Both $D_{1}$ and $D_{2}$ are randomly selected. We conduct two experiments: one is the comparison between the ADL and some existing methods to show the advantage of the ADL in terms of time complexity. Another is used to indicate the low space complexity of the ADL.
In our first experiment, we choose $n=200,p=500$ and $s_{0}=6$ and compare the performance of the ADL with offline debiased lasso (deLasso) in van de Geer et al., (2014), link-specific weighting (LSW) method proposed in Cai et al., (2023) and online debiased lasso (ODL) proposed in Luo et al., 2023a . In addition, we check the robustness of these methods with respect to violating the bounded individual probability condition or not. For this reason, we consider two types of covariance matrix $\bm{\Sigma}=0.1\times\{0.5^{|i-j|}\}_{i,j=1,\dots,p}$ and $\bm{\Sigma}=\{0.5^{|i-j|}\}_{i,j=1,\dots,p}$. This setting is similar to the setting in Cai et al., (2023). The former will ensure the bounded individual probability condition holds, while the latter will break this condition since the resulting probability can be pretty close to 00. Besides that, the consideration of other covariance matrices, such as the block-wise Toeplitz matrix, as well as the choice of the hyper-parameters is given in the supplementary material.
The performance is based on four criteria: (1) Absolute bias: the absolute distance between estimator and ground truth. (2) Converge probability of 95% confidence interval (3) Length of 95% confidence interval and (4) the computation times. Note that there are three different categories among $\bm{\beta}^{*}_{i},i\in[p]$, that is $0,1$ and $-1$. We choose three elements from each category and calculate the corresponding debiased estimator respectively. The averaged performances for each category are recorded. The experiment is conducted with 200 replications and the results are summarized in Tables 2 and 3.
Table 2 presents the results when the covariance matrix $\bm{\Sigma}=0.1\times\{0.5^{|i-j|}\}_{i,j=1,\dots,p}$. Notably, all four methods are able to achieve the desired coverage probability. Our ADL estimator takes the minimum time and can still achieve the comparable length of the confidence interval. In Table 3, when the bounded individual probability condition is violated, both deLasso and its online variant ODL fail to achieve $95\%$ coverage probability. This phenomenon has also been observed in Cai et al., (2023). In contrast, LSW and ADL are still able to achieve $95\%$ coverage probability. This result further supports our finding that similar to LSW, the ADL does not depend on the bounded individual probability condition. Compared with LSW, the ADL method is around 50 to 60 times faster with similar statistical efficiency. In summary, the ADL enjoys a small time complexity and still yields high statistical accuracy.
In the second setting, we aim to demonstrate the low space complexity of the ADL. We increase the value of $p$ to $20000$ and choose $n=1000,s_{0}=20$ with $\bm{\Sigma}=\{0.5^{|i-j|}\}_{i,j=1,\dots,p}$. Under this circumstance, the computation of deLasso, LSW, or ODL is out of memory and only ADL is computable. Similar to the first setting, we construct the confidence intervals for three randomly selected parameters from each category of $\bm{\beta}^{*}$. The averaged result in each category is summarized in Figure 3. In Figure 3, the lines and shadow stand for the change of debiased estimator and the length of 95% confidence interval along with the increase of data size respectively. We can observe that the estimator is closed to the corresponding true value in every category, which shows the accuracy of our method. Another unexpected observation is that the length of the confidence interval decreases as the number of data increases. The coverage probability also achieves 95% for each case and the detailed results can be found in the supplementary material. The second experiment shows the ADL has low space complexity but still maintains the desired statistical accuracy.
Table 2: Simulation results over 200 replications with $n=200$, $p=500$, $s_{0}=6$, $\bm{\Sigma}=0.1\times\{0.5^{|i-j|}\}_{i,j=1,\dots,p}$.
$\bm{\beta}_{k}^{*}$ deLasso LSW ODL ADL
Sample size $n$ 200 200 80 140 200 80 140 200
Absolute bias 0 0.410 0.368 0.066 0.043 0.038 0.716 0.510 0.435
1 0.436 0.532 0.259 0.219 0.181 0.707 0.486 0.417
-1 0.411 0.521 0.231 0.204 0.178 0.707 0.545 0.436
Coverage probability 0 0.953 0.988 0.971 0.963 0.958 0.968 0.963 0.950
1 0.930 0.925 0.963 0.951 0.946 0.963 0.975 0.961
-1 0.952 0.945 0.975 0.958 0.953 0.978 0.956 0.950
Confidence interval length 0 2.036 2.260 4.061 2.297 1.771 3.903 2.665 2.143
1 2.036 2.265 4.059 2.293 1.766 3.916 2.665 2.151
-1 2.045 2.284 4.09 2.304 1.776 3.855 2.654 2.140
Time $(s)$ 7.214 20.22 0.539 0.324
Table 3: Simulation results averaged over 200 replications with $n=200$, $p=500$, $s_{0}=6$, $\bm{\Sigma}=\{0.5^{|i-j|}\}_{i,j=1,\dots,p}$.
$\bm{\beta}_{k}^{*}$ deLasso LSW ODL ADL
Sample size $n$ 200 200 80 140 200 80 140 200
Absolute bias 0 0.114 0.113 0.018 0.015 0.013 0.509 0.344 0.295
1 0.325 0.260 0.556 0.505 0.488 0.520 0.369 0.302
-1 0.326 0.259 0.574 0.522 0.481 0.539 0.384 0.316
Coverage probability 0 0.983 0.996 1.000 0.999 0.997 0.948 0.965 0.956
1 0.598 0.936 0.956 0.790 0.646 0.928 0.940 0.953
-1 0.610 0.946 0.958 0.828 0.676 0.918 0.926 0.928
Confidence interval length 0 0.753 1.207 1.834 1.354 1.133 2.487 1.751 1.436
1 0.769 1.261 1.833 1.355 1.131 2.546 1.784 1.458
-1 0.767 1.271 1.828 1.348 1.130 2.496 1.759 1.449
Time $(s)$ 6.550 20.04 0.628 0.365
Figure 3: Simulation results averaged over 200 replications (each takes around 53.54s) with $n=1000$, $p=20000$, $s_{0}=20$, $\bm{\Sigma}=\{0.5^{|i-j|}\}_{i,j=1,\dots,p}$.
4.2 Real data example: Identifying malicious URLs
With the rapid development of online technology, it has become more convenient for people to communicate with each other. Nowadays, human communication is not only limited to emails but also through websites and social networking platforms. However, this has also created opportunities for criminals to take advantage of the situation. Every day we still receive lots of spam emails, malicious links, and phishing websites from the Internet. How to identify fake information is important, which can help us to avoid the leakage of personal information and the loss of property. For this purpose, Ma et al., (2009) considered an online algorithm to predict malicious URLs. The data is collected in 120 days, resulting in 2.4 million data and around 3.2 million features. The authors compared several methods and reported the cumulative prediction errors. Specifically, they considered the logistics regression with the stochastic gradient descent, whose prediction error is below 5%.
Different from the prediction error, we are more interested in which attributes are significant to make the malicious URL distinguished from the benign one. We take one further step to conduct inference with the approximated debiased lasso algorithm. We evaluate our algorithm in two aspects: accuracy and computational efficiency. However, it is pretty hard to judge accuracy since all features are anonymized. For this reason, we add one fake feature, which is randomly generated from the standard normal distribution, as a benchmark. In addition to the fake feature, we also conduct inference for features 16 and 32. The results of the first-day data (with the size 20,000) are presented in Figure 4.
Similar to Figure 3, the lines and shadow show the ADL estimator and the 95% confidence interval respectively. According to the right panel, we observe that the confidence interval of the fake feature always contains zero, which indicates the accuracy of ADL. Besides that, the results from the left and middle panels imply that Feature 16 is significant while Feature 32 is not under the significance level 0.05.
In terms of computational efficiency, the online algorithm should be able to give the response for the current data prior to the arrival of the next data. The data in Ma et al., (2009) is collected in an online fashion, with approximately $4.32s$ for each. In contrast, the ADL costs $0.3s$ for processing each single data. This result reveals the efficiency of the ADL and its ability to conduct online inference with ultra-high-dimensional data.
Figure 4: 95% confidence interval estimate for Feature 16 (left), Feature 32 (middle), and Fake Feature (right).
5 Conclusion
High-dimensional inference has been well investigated in the last decade. However, it still lacks an efficient method to conduct the inference in a real online manner. In this work, we contribute to the literature by proposing the ADL to deal with the high-dimensional generalized linear model. While maintaining comparable accuracy to the offline method, the ADL requires much less time and space complexities.
Several directions deserve further investigation. First, the ADL is designed for an objective function with a lasso penalty. It is unclear what if some other penalty functions, such as SCAD (Fan and Li,, 2001) and MCP (Zhang,, 2010), are used. Second, our work focuses on the parametric models. It is interesting to investigate whether our approach can be extended to the model with a non-parametric component to conduct semi-parametric inference. Last but not least, this work assumes that sequentially collected data points are independent and identically distributed (i.i.d). How to deal with non-i.i.d. design, such as the covariate shift, under an online setting is still an open question.
Appendix A Proof related to the Convergence
A.1 Hyper-parameters used in proof
We first list the hyper-parameters. The constants $c_{i},i\in[4]$ are universal constants, whose exact value will be given in the supplementary material. Recall that $d_{0}=\lVert\bm{\beta}^{(0)}-\bm{\beta}^{*}\lVert_{1}$. Let $\omega_{1}=9\sqrt{\log\log p}$ and $R_{1}=c_{1}d_{0}$. Since we always take the zero point as the initial value, $R_{1}=c_{1}\lVert\bm{\beta}^{*}\lVert_{1}$. Moreover,
$R_{k}=R_{1}/(\sqrt{2})^{k-1},\ \xi_{k}=\kappa_{1}^{2}M_{\Phi}R_{k},\ \zeta_{k}^{2}=2\{\kappa_{2}\log(p)+M_{\Phi}^{2}R_{k}^{2}\log^{2}(p)\},$ (20)
and $\omega_{k}^{2}=\omega_{1}^{2}+24\log(k)$ for $k\in[K]$. Based on these hyper-parameters, we further define $T_{0}=1$ and
$T_{k}=(2T_{k-1})\vee\left[c_{2}\frac{s_{0}^{2}}{R_{k}^{2}}\big{\{}(\xi_{k}^{2}+\zeta_{k}^{2})\log p+\omega_{k}^{2}\zeta_{k}^{2}\big{\}}+c_{2}\log p\right],$
$\alpha_{k}=c_{3}\sqrt{\frac{\log p}{R_{k}^{2}(\xi_{k}^{2}+\lambda_{k}^{2}+\zeta_{k}^{2})}},\ \lambda_{k}^{2}=\frac{c_{4}R_{k}}{s_{0}\sqrt{T_{k}}}\Big{(}{\sqrt{(\xi_{k}^{2}+\zeta_{k}^{2})\log p}}+{\zeta_{k}\omega_{k}}\Big{)}.$
A.2 Lemmas
Let $\bar{F}(\bm{\beta};\lambda_{k}):=\mathbb{E}\{F_{n}(\bm{\beta};\lambda_{k})\}$ and
$\bm{\beta}^{(n_{k})}=\operatorname*{arg\,min\;}_{\bm{\beta}\in\mathbb{R}^{n}}\bar{F}(\bm{\beta};\lambda_{k}).$ (21)
We first show the convergence rate within a single epoch by the following lemma:
Lemma A.1.
Suppose that Assumptions 1,2,3 hold and $\lVert\widehat{\bm{\beta}}^{(n_{k-1})}-{\bm{\beta}}^{(n_{k})}\lVert_{1}\leq R_{k}.$ Then,
$\lVert\widehat{\bm{\beta}}^{(n_{k})}-{\bm{\beta}}^{*}\lVert_{1}\leq R_{k+1},\ \lVert\widehat{\bm{\beta}}^{(n_{k})}-{\bm{\beta}}^{(n_{k+1})}\lVert_{1}\leq R_{k+1},\ \lVert\widehat{\bm{\beta}}^{(n_{k})}-{\bm{\beta}}^{*}\lVert_{2}\leq R_{k+1}/\sqrt{s_{0}},$
holds with probability at least $1-4\exp(-\omega_{k}^{2}/12)$.
Proof of Theorem 3.1.
Note that we choose $\widehat{\bm{\beta}}^{(0)}$ as the zero point. Since $\lVert\widehat{\bm{\beta}}^{(0)}-{\bm{\beta}}^{*}\lVert_{1}\leq R_{1}$ and $\lVert{\bm{\beta}}^{*}\lVert_{1}\geq\lVert{\bm{\beta}}^{(n_{1})}\lVert_{1}$, we obtain $\lVert\widehat{\bm{\beta}}^{(0)}-{\bm{\beta}}^{(n_{1})}\lVert_{1}\leq R_{1}$. Then, according to Lemma A.1, we obtain that
$\lVert\widehat{\bm{\beta}}^{(n_{k})}-\bm{\beta}^{*}\lVert_{1}\leq R_{k+1}^{2}=R_{1}^{2}/2^{k},\ k\in[K],$
holds uniformly with probability at least $1-4\sum_{i=1}^{K}\exp(-\omega_{i}^{2}/12).$ Based on the choice of $\omega_{i}$, we have $4\sum_{i=1}^{k}\exp(-\omega_{i}^{2}/12)\leq 7(\log p)^{-6}.$ Meanwhile,
$n_{k}$ $=\sum_{i=1}^{k}T_{i}=\sum_{i=1}^{k}\left\{(2T_{i-1})\vee\left[c_{2}\frac{s_{0}^{2}}{R_{i}^{2}}\big{\{}(\xi_{i}^{2}+\zeta_{i}^{2})\log(p)+(\omega_{1}^{2}+24\log(i))\zeta_{i}^{2}\big{\}}+c_{2}\log(p)\right]\right\}$
$\leq\sum_{i=1}^{k}\left\{(2T_{i-1})\vee\left[c_{2}\frac{s_{0}^{2}}{R_{i}^{2}}\big{\{}(\xi_{i}^{2}+\zeta_{i}^{2})\log(p)+(\omega_{1}^{2}+24\log(k))\zeta_{i}^{2}\big{\}}+c_{2}\log(p)\right]\right\}$
$\leq\sum_{i=1}^{k}\left\{(2T_{i-1})\vee\left[2c_{2}\frac{s_{0}^{2}}{R_{i-1}^{2}}\big{\{}(\xi_{i}^{2}+\zeta_{i}^{2})\log(p)+(\omega_{1}^{2}+24\log(k))\zeta_{i}^{2}\big{\}}+2c_{2}\log(p)\right]\right\}.$
Note that $\xi_{i+1}\leq\xi_{i}$ and $\zeta_{i+1}\leq\zeta_{i}$ for $i\in[K-1]$. Then, we obtain
$n_{k}$ $\leq\sum_{i=1}^{k}\left\{(2T_{i-1})\vee\left[2c_{2}\frac{s_{0}^{2}}{R_{i-1}^{2}}\big{\{}(\xi_{i-1}^{2}+\zeta_{i-1}^{2})\log(p)+(\omega_{1}^{2}+24\log(k))\zeta_{i-1}^{2}\big{\}}+2c_{2}\log(p)\right]\right\}$
$=\sum_{i=1}^{k}2c_{2}\left[\frac{s_{0}^{2}}{R_{i-1}^{2}}\big{\{}(\xi_{i-1}^{2}+\zeta_{i-1}^{2})\log(p)+(\omega_{1}^{2}+24\log(k))\zeta_{i-1}^{2}\big{\}}+\log(p)\right]$
$\leq\sum_{i=1}^{k}2^{i-1}c_{2}\left[\frac{s_{0}^{2}}{R_{1}^{2}}\big{\{}(\xi_{1}^{2}+\zeta_{1}^{2})\log(p)+(\omega_{1}^{2}+24\log(k))\zeta_{1}^{2}\big{\}}+\log(p)\right]$
$\leq 2^{k}\left(T_{1}+\frac{24c_{2}\zeta_{1}^{2}s_{0}^{2}\log(k)}{R_{1}^{2}}\right).$
Recall that $R_{1}\gtrsim 1$ and $p\geq n$. Then, there exists a universal constant $c_{5}$ such that
$T_{1}\leq c_{5}s_{0}^{2}(\log p)^{3},\ \frac{24\zeta_{1}^{2}s_{0}^{2}\gamma^{2}\log k}{R_{1}^{2}}\leq c_{5}s_{0}^{2}(\log p)^{3}.$
Consequently,
$\lVert\widehat{\bm{\beta}}^{(n_{k})}-\bm{\beta}^{*}\lVert_{1}\leq c_{6}s_{0}R_{1}\sqrt{\frac{(\log p)^{3}}{n_{k}}},\ k\in[L_{T}],$
holds uniformly with probability at least $1-7(\log p)^{-6}.$ Since $2T_{k}\leq T_{k+1}$ for $k\in[K-1]$, we let $c_{7}=\sqrt{2}c_{6}$ and draw the conclusion that
$\lVert\widehat{\bm{\beta}}^{(i)}-\bm{\beta}^{*}\lVert_{1}\leq c_{7}s_{0}R_{1}\sqrt{\frac{(\log p)^{3}}{i}},i\in[n],$
holds uniformly with probability at least $1-7(\log p)^{-6}.$ We finish the proof of Theorem 3.1. ∎
Appendix B Convergence rate in nodewise lasso
Recall that $d_{1}=\lVert\widehat{\bm{\gamma}}_{\cdot,j}^{(0)}-\bm{\gamma}_{\cdot,j}^{*}\lVert_{1}.$ Let $\omega_{1}^{\prime}=9\sqrt{\log\log p}$, $c_{i}^{\prime},i\in[4]$ denote some universal constants, and
$R_{1}^{\prime}=c_{1}^{\prime}\left[d_{1}+R_{1}+\{(s_{0}R_{1})/(s_{1})\}\right].$
In addition,
$R_{k}^{\prime}=R_{1}^{\prime}/(\sqrt{2})^{k-1},\ \xi_{k}^{\prime}=\kappa_{1}^{2}M_{\Phi}R_{k}^{\prime},\ (\zeta_{k}^{\prime})^{2}=2\{\kappa_{1}M_{2}^{2}\log(p)+M_{\Phi}^{2}R_{k}^{\prime 2}\log^{2}(p)\},$ (22)
and $(\omega_{k}^{\prime})^{2}=(\omega_{1}^{\prime})^{2}+24\log(k)$ for $k\in[K^{\prime}]$. With these hyper-parameters, we further define $T^{\prime}_{0}=1$ and
$T_{k}^{\prime}=(2T_{k-1}^{\prime})\vee\left[c_{2}^{\prime}\frac{s_{1}^{2}}{R_{k}^{\prime 2}}\{(\xi_{k}^{\prime 2}+\zeta_{k}^{\prime 2})\log p+\omega_{k}^{\prime 2}\zeta_{k}^{\prime 2}\}+c_{2}^{\prime}\log p\right],$
$\alpha_{k}^{\prime}=c_{3}^{\prime}\sqrt{\frac{\log p}{R_{k}^{\prime 2}(\xi_{k}^{\prime 2}+\lambda_{k}^{\prime 2}+\zeta_{k}^{\prime 2})}},\ (\lambda_{k}^{\prime})^{2}=\frac{c_{4}^{\prime}R_{k}^{\prime}}{s_{1}\sqrt{T_{k}^{\prime}}}\Big{(}\sqrt{(\xi_{k}^{\prime 2}+\zeta_{k}^{\prime 2})\log p}+\zeta_{k}^{\prime}\omega_{k}^{\prime}\Big{)}.$
We also provide the convergence rate of the nodewise lasso using the adaptive RADAR.
Lemma B.1.
Suppose that Assumptions 1–4 and (17) hold. If $\lVert{\widehat{\bm{\gamma}}_{\cdot,j}}^{(n_{k-1}^{\prime})}-\bm{\gamma}_{\cdot,j}^{(n_{k}^{\prime})}\lVert_{1}\leq R_{k}^{\prime}$, then
$\lVert{\widehat{\bm{\gamma}}_{\cdot,j}}^{(n_{k}^{\prime})}-\bm{\gamma}_{\cdot,j}^{(n_{k}^{\prime})}\lVert_{1}\leq R_{k+1}^{\prime},\ \lVert{\widehat{\bm{\gamma}}_{\cdot,j}}^{(n_{k}^{\prime})}-\bm{\gamma}_{\cdot,j}^{*}\lVert_{1}\leq R_{k+1}^{\prime}$ (23)
holds with probability at least $1-4\exp(-\omega_{k}^{\prime 2}/12)$.
Proof of Theorem 3.2.
[…]
$\lVert{\widehat{\bm{\gamma}}_{\cdot,j}}^{(0)}-\bm{\gamma}_{\cdot,j}^{(n_{1}^{\prime})}\lVert_{1}$ $\leq\lVert{\widehat{\bm{\gamma}}_{\cdot,j}}^{(0)}-\bm{\gamma}_{\cdot,j}^{*}\lVert_{1}+\lVert\bm{\gamma}_{\cdot,j}^{*}-\bm{\gamma}_{\cdot,j}^{(n_{1}^{\prime})}\lVert_{1}\leq\lVert{\widehat{\bm{\gamma}}_{\cdot,j}}^{(0)}-\bm{\gamma}_{\cdot,j}^{*}\lVert_{1}+CR_{1}/\sqrt{2}\leq R_{1}^{\prime}.$
Then, by Lemma B.1,
$\lVert\widehat{\bm{\gamma}}_{\cdot,j}^{(n_{k}^{\prime})}-\bm{\gamma}_{\cdot,j}^{*}\lVert_{1}\leq R_{k+1}^{\prime}=R_{1}^{\prime}/(\sqrt{2})^{k},\ k\in[K],$
holds uniformly with probability at least $1-4\sum_{i=1}^{K}\exp(-\omega_{i}^{\prime 2}/12).$ Based on the choice of $\omega_{i}^{\prime}$, we have $4\sum_{i=1}^{k}\exp(-\omega_{i}^{\prime 2}/12)\leq 7(\log p)^{-6}.$ Since
$(n_{k}^{\prime}-n_{1})=\sum_{i=1}^{k}T_{i}^{\prime},\ \xi^{\prime}_{i}\leq\xi^{\prime}_{i-1},\text{ and }\zeta^{\prime}_{i}\leq\zeta^{\prime}_{i-1}$
holds for $i\in[K^{\prime}]$, we follow the same procedure as the proof of Theorem 3.1. Consequently,
$\lVert\widehat{\bm{\gamma}}_{\cdot,j}^{(n_{k}^{\prime})}-\bm{\gamma}_{\cdot,j}^{*}\lVert_{1}\leq c_{6}^{\prime}s_{1}R_{1}^{\prime}\sqrt{\frac{(\log p)^{3}}{n_{k}^{\prime}-n_{1}}},\ k\in[L_{T}],$
holds uniformly with probability at least $1-7(\log p)^{-6}.$ Since $2T_{k}^{\prime}\leq T_{k+1}^{\prime}$ for $k\in[K^{\prime}-1]$, we let $c_{7}^{\prime}=\sqrt{2}c_{6}^{\prime}$ and obtain that
$\lVert\widehat{\bm{\gamma}}_{\cdot,j}^{(i)}-\bm{\gamma}_{\cdot,j}^{*}\lVert_{1}\leq c_{7}s_{1}R_{1}^{\prime}\sqrt{\frac{(\log p)^{3}}{i-n_{1}}},i=n_{1}^{\prime},n_{1}^{\prime}+1,\ldots,n,$
[…]
$\lVert\widehat{\bm{\gamma}}_{\cdot,j}^{(i)}-\bm{\gamma}_{\cdot,j}^{*}\lVert_{1}\leq c_{7}s_{1}R_{1}^{\prime}\sqrt{\frac{(\log p)^{3}}{i-n_{1}}},i=n_{1}^{\prime},n_{1}^{\prime}+1,\ldots,n,$
holds uniformly with probability at least $1-14(\log p)^{-6}.$
Appendix C Proof of Theorem 3.3
Proof.
We first state some facts about the length of the epoch, that is,
[…]
Recall that
${\widehat{\bm{\beta}}}^{(n)}_{j,\text{de}}=\widehat{{\bm{\beta}}}^{(n)}_{j}-\frac{\{\bm{a}^{(n)}_{1}\}^{\intercal}{\widehat{\bm{\gamma}}_{\cdot,j}}^{(n)}+\{\bm{a}^{(n)}_{2}\}^{\intercal}\widehat{{\bm{\beta}}}^{(n)}-a_{3}^{(n)}}{{a}_{4}^{(n)}},$
where $\bm{a}_{1}^{(n)},\bm{a}_{2}^{(n)},a_{3}^{(n)}$ and $a_{4}^{(n)}$ are defined in (12). We have
${a}_{4}^{(n)}({\widehat{\bm{\beta}}}^{(n)}_{j,\text{de}}-{{\bm{\beta}}}^{*}_{j})={a}_{4}^{(n)}(\widehat{{\bm{\beta}}}^{(n)}_{j}-{{\bm{\beta}}}^{*}_{j})-\{\bm{a}^{(n)}_{1}\}^{\intercal}{\widehat{\bm{\gamma}}_{\cdot,j}}^{(n)}-\{\bm{a}^{(n)}_{2}\}^{\intercal}\widehat{{\bm{\beta}}}^{(n)}+a_{3}^{(n)}.$
Since we split $n$ data into $T_{K}$ epochs, we can rewrite $\bm{a}_{1}^{(n)}$ with (4) as
$\bm{a}_{1}^{(n)}=\sum_{k=l+1}^{K}\sum_{i=n_{k-1}+1}^{n_{k}}\bm{x}_{i}\{-y_{i}+\dot{\Phi}(\bm{x}^{\intercal}_{i}{\widehat{\bm{\beta}}^{(n_{k-1})}})\}.$
Let ${\bm{m}}_{1}^{(k)}=\sum_{i=n_{k-1}+1}^{n_{k}}\bm{x}_{i}\left\{-y_{i}+\dot{\Phi}(\bm{x}^{\intercal}_{i}\widehat{\bm{\beta}}^{(n_{k-1})})\right\}$. It follows that
${\bm{m}}_{1}^{(k)}$ $=\sum_{i=n_{k-1}+1}^{n_{k}}\bm{x}_{i}\left\{-y_{i}+\dot{\Phi}(\bm{x}^{\intercal}_{i}{\bm{\beta}}^{*})\right\}-\sum_{i=n_{k-1}+1}^{n_{k}}\bm{x}_{i}\left\{\dot{\Phi}(\bm{x}^{\intercal}_{i}{\bm{\beta}}^{*})-\dot{\Phi}(\bm{x}^{\intercal}_{i}\widehat{\bm{\beta}}^{(n_{k-1})})\right\}$
$=$ $\sum_{i=n_{k-1}+1}^{n_{k}}\bm{x}_{i}\left\{-y_{i}+\dot{\Phi}(\bm{x}^{\intercal}_{i}{\bm{\beta}}^{*})\right\}-\sum_{i=n_{k-1}+1}^{n_{k}}\bm{x}_{i}\ddot{\Phi}(\eta_{k,i})\bm{x}^{\intercal}_{i}({\bm{\beta}}^{*}-\widehat{\bm{\beta}}^{(n_{k-1})})$
$=$ $\sum_{i=n_{k-1}+1}^{n_{k}}\bm{x}_{i}\left\{-y_{i}+\dot{\Phi}(\bm{x}^{\intercal}_{i}{\bm{\beta}}^{*})\right\}-\sum_{i=n_{k-1}+1}^{n_{k}}\bm{x}_{i}\ddot{\Phi}(\eta_{k,i})\bm{x}^{\intercal}_{i}({\bm{\beta}}^{*}-\widehat{\bm{\beta}}^{(n_{k-1})})$
$=$ $\sum_{i=n_{k-1}+1}^{n_{k}}\bm{x}_{i}\left\{-y_{i}+\dot{\Phi}(\bm{x}^{\intercal}_{i}{\bm{\beta}}^{*})\right\}-\sum_{i=n_{k-1}+1}^{n_{k}}\bm{x}_{i}\ddot{\Phi}(\bm{x}^{\intercal}_{i}\widehat{\bm{\beta}}^{(n_{k-1})})\bm{x}^{\intercal}_{i}({\bm{\beta}}^{*}-\widehat{\bm{\beta}}^{(n_{k-1})})$ (25)
$-\sum_{i=n_{k-1}+1}^{n_{k}}\bm{x}_{i}\left\{\ddot{\Phi}(\eta_{k,i})-\ddot{\Phi}(\bm{x}^{\intercal}_{i}\widehat{\bm{\beta}}^{(n_{k-1})})\right\}\bm{x}^{\intercal}_{i}({\bm{\beta}}^{*}-\widehat{\bm{\beta}}^{(n_{k-1})}),$
where $\eta_{k,i}\in\big{(}\bm{x}^{\intercal}_{i}{\bm{\beta}}^{*},\bm{x}^{\intercal}_{i}\widehat{\bm{\beta}}^{(n_{k-1})}\big{)}.$ We further obtain that
$\bm{a}_{1}^{(n)}$ $=\sum_{i=n_{l}+1}^{n}\bm{x}_{i}\left\{-y_{i}+\dot{\Phi}(\bm{x}^{\intercal}_{i}{\bm{\beta}}^{*})\right\}-\sum_{k=l+1}^{K}\sum_{i=n_{k-1}+1}^{n_{k}}\bm{x}_{i}\ddot{\Phi}(\bm{x}^{\intercal}_{i}\widehat{\bm{\beta}}^{(n_{k-1})})\bm{x}^{\intercal}_{i}({\bm{\beta}}^{*}-\widehat{\bm{\beta}}^{(n_{k-1})})$
$-\sum_{k=l+1}^{K}\sum_{i=n_{k-1}+1}^{n_{k}}\bm{x}_{i}\left\{\ddot{\Phi}(\eta_{k,i})-\ddot{\Phi}(\bm{x}^{\intercal}_{i}\widehat{\bm{\beta}}^{(n_{k-1})})\right\}\bm{x}^{\intercal}_{i}({\bm{\beta}}^{*}-\widehat{\bm{\beta}}^{(n_{k-1})}).$
Meanwhile,
$\{\bm{a}^{(n)}_{2}\}^{\intercal}\widehat{{\bm{\beta}}}^{(n)}-a_{3}^{(n)}=\sum_{k=l+1}^{K}\sum_{i=n_{k-1}+1}^{n_{k}}\bm{x}_{i}^{\intercal}{\widehat{\bm{\gamma}}_{\cdot,j}}^{(i)}\ddot{\Phi}(\bm{x}^{\intercal}_{i}\widehat{\bm{\beta}}^{(n_{k-1})})\bm{x}^{\intercal}_{i}(\widehat{\bm{\beta}}^{(n)}-\widehat{\bm{\beta}}^{(n_{k-1})}).$
As a result,
${a}_{4}^{(n)}({\widehat{\bm{\beta}}}^{(n)}_{j,\text{de}}-{{\bm{\beta}}}^{*}_{j})=-\{\bm{\gamma}_{\cdot,j}^{*}\}^{\intercal}\sum_{i=n_{l}+1}^{n}\bm{x}_{i}\left\{-y_{i}+\dot{\Phi}(\bm{x}^{\intercal}_{i}{\bm{\beta}}^{*})\right\}+\Delta_{1}+\Delta_{2}+\Delta_{3}+\Delta_{4},$
where
$\Delta_{1}$ $=\sum_{k=l+1}^{K}\sum_{i=n_{k-1}+1}^{n_{k}}\{{\widehat{\bm{\gamma}}_{\cdot,j}}^{(n)}-{\widehat{\bm{\gamma}}_{\cdot,j}}^{(i)}\}^{\intercal}\bm{x}_{i}\ddot{\Phi}(\bm{x}^{\intercal}_{i}\widehat{\bm{\beta}}^{(n_{k-1})})\bm{x}^{\intercal}_{i}({\bm{\beta}}^{*}-\widehat{\bm{\beta}}^{(n_{k-1})}),$
$\Delta_{2}$ $=\sum_{k=l+1}^{K}\sum_{i=n_{k-1}+1}^{n_{k}}\{{\widehat{\bm{\gamma}}_{\cdot,j}}^{(i)}\}^{\intercal}\bm{x}_{i}\ddot{\Phi}(\bm{x}^{\intercal}_{i}\widehat{\bm{\beta}}^{(n_{k-1})})\left\{\bm{x}^{\intercal}_{i}-(\bm{x}_{i})_{j}\bm{e}_{j}^{\intercal}\right\}({\bm{\beta}}^{*}-\widehat{\bm{\beta}}^{(n)}),$
$\Delta_{3}$ =$=\{{\widehat{\bm{\gamma}}_{\cdot,j}}^{(n)}\}^{\intercal}\sum_{k=l+1}^{K}\sum_{i=n_{k-1}+1}^{n_{k}}\bm{x}_{i}\left\{\ddot{\Phi}(\eta_{k,i})-\ddot{\Phi}(\bm{x}^{\intercal}_{i}\widehat{\bm{\beta}}^{(n_{k-1})})\right\}\bm{x}^{\intercal}_{i}({\bm{\beta}}^{*}-\widehat{\bm{\beta}}^{(n_{k-1})}),$
$\Delta_{4}$ =$=\{\bm{\gamma}_{\cdot,j}^{*}-{\widehat{\bm{\gamma}}_{\cdot,j}}^{(n)}\}^{\intercal}\sum_{i=n_{l}+1}^{n}\bm{x}_{i}\left\{-y_{i}+\dot{\Phi}(\bm{x}^{\intercal}_{i}{\bm{\beta}}^{*})\right\}.$
Lemmas C.1C.4 show that $|\Delta_{i}|=o_{\mathbb{P}}(\sqrt{n}),i=1,2,3,4.$ Consequently,
${a}_{4}^{(n)}({\widehat{\bm{\beta}}}^{(n)}_{j,\text{de}}-{{\bm{\beta}}}^{*}_{j})=-\{\bm{\gamma}_{\cdot,j}^{*}\}^{\intercal}\sum_{i=n_{l}+1}^{n}\bm{x}_{i}\left\{-y_{i}+\dot{\Phi}(\bm{x}^{\intercal}_{i}{\bm{\beta}}^{*})\right\}+o_{\mathbb{P}}(\sqrt{n}).$
Based on the central limit theorem,
$\frac{1}{\nu_{j}^{(n)}}({\widehat{\bm{\beta}}}^{(n)}_{j,\text{de}}-{{\bm{\beta}}}^{*}_{j})\to\mathcal{N}(0,1)$
in distribution as $n\to\infty$, where
$\nu_{j}^{(n)}=\frac{\sqrt{n}}{|{a}_{4}^{(n)}|}\sqrt{\mathbb{E}\left[(\bm{x}^{\intercal}\bm{\gamma}_{\cdot,j}^{*})\{-y+\dot{\Phi}(\bm{x}^{\intercal}{\bm{\beta}}^{*})\}\right]^{2}}.$
Then, by Slutsky’s theorem, the proof is complete as long as
$\frac{1}{n-n_{l}}\sum_{i=n_{l}+1}^{n}(\bm{x}_{i}^{\intercal}{\widehat{\bm{\gamma}}_{\cdot,j}}^{(i)})^{2}\left\{\dot{\Phi}(\bm{x}^{\intercal}_{i}\widehat{\bm{\beta}}^{(i)})-y_{i}\right\}^{2}\to\mathbb{E}\left[(\bm{x}^{\intercal}\bm{\gamma}_{\cdot,j}^{*})\{-y+\dot{\Phi}(\bm{x}^{\intercal}{\bm{\beta}}^{*})\}\right]^{2},$ (26)
as $n\to\infty.$ The remaining part is to prove (26). By the law of large numbers, it suffices to show that
$\sum_{i=n_{l}+1}^{n}(\bm{x}_{i}^{\intercal}{\widehat{\bm{\gamma}}_{\cdot,j}}^{(i)})^{2}\left\{\dot{\Phi}(\bm{x}^{\intercal}_{i}\widehat{\bm{\beta}}^{(i)})-y_{i}\right\}^{2}-\sum_{i=n_{l}+1}^{n}(\bm{x}_{i}^{\intercal}\bm{\gamma}_{\cdot,j}^{*})^{2}\left\{\dot{\Phi}(\bm{x}^{\intercal}_{i}{\bm{\beta}}^{*})-y_{i}\right\}^{2}=\mathcal{O}_{\mathbb{P}}(n).$ (27)
We first write the left-hand side of (27) into the following two parts:
$\sum_{i=n_{l}+1}^{n}(\bm{x}_{i}^{\intercal}{\widehat{\bm{\gamma}}_{\cdot,j}}^{(i)})^{2}\left\{\dot{\Phi}(\bm{x}^{\intercal}_{i}\widehat{\bm{\beta}}^{(i)})-y_{i}\right\}^{2}-\sum_{i=n_{l}+1}^{n}(\bm{x}_{i}^{\intercal}\bm{\gamma}_{\cdot,j}^{*})^{2}\left\{\dot{\Phi}(\bm{x}^{\intercal}_{i}{\bm{\beta}}^{*})-y_{i}\right\}^{2}$
$=$ $\sum_{k=l+1}^{K}\sum_{i=n_{k-1}+1}^{n_{k}}\left[(\bm{x}_{i}^{\intercal}{\widehat{\bm{\gamma}}_{\cdot,j}}^{(i)})^{2}\left\{\dot{\Phi}(\bm{x}^{\intercal}_{i}\widehat{\bm{\beta}}^{(n_{k-1})})-y_{i}\right\}^{2}-(\bm{x}_{i}^{\intercal}{\widehat{\bm{\gamma}}_{\cdot,j}}^{(i)})^{2}\left\{\dot{\Phi}(\bm{x}^{\intercal}_{i}{\bm{\beta}}^{*})-y_{i}\right\}^{2}\right]$
$+\sum_{k=l+1}^{K}\sum_{i=n_{k-1}+1}^{n_{k}}\left[(\bm{x}_{i}^{\intercal}{\widehat{\bm{\gamma}}_{\cdot,j}}^{(i)})^{2}\left\{\dot{\Phi}(\bm{x}^{\intercal}_{i}{\bm{\beta}}^{*})-y_{i}\right\}^{2}-(\bm{x}_{i}^{\intercal}\bm{\gamma}_{\cdot,j}^{*})^{2}\left\{\dot{\Phi}(\bm{x}^{\intercal}_{i}{\bm{\beta}}^{*})-y_{i}\right\}^{2}\right].$
For the first part,
$\sum_{k=l+1}^{K}\sum_{i=n_{k-1}+1}^{n_{k}}\left[(\bm{x}_{i}^{\intercal}{\widehat{\bm{\gamma}}_{\cdot,j}}^{(i)})^{2}\left\{\dot{\Phi}(\bm{x}^{\intercal}_{i}\widehat{\bm{\beta}}^{(n_{k-1})})-y_{i}\right\}^{2}-(\bm{x}_{i}^{\intercal}{\widehat{\bm{\gamma}}_{\cdot,j}}^{(i)})^{2}\left\{\dot{\Phi}(\bm{x}^{\intercal}_{i}{\bm{\beta}}^{*})-y_{i}\right\}^{2}\right]$ (28)
$=$ $\sum_{k=l+1}^{K}\sum_{i=n_{k-1}+1}^{n_{k}}\left[(\bm{x}_{i}^{\intercal}{\widehat{\bm{\gamma}}_{\cdot,j}}^{(i)})^{2}\left\{\dot{\Phi}(\bm{x}^{\intercal}_{i}\widehat{\bm{\beta}}^{(n_{k-1})})+\dot{\Phi}(\bm{x}^{\intercal}_{i}{\bm{\beta}}^{*})-2y_{i}\right\}\left\{\dot{\Phi}(\bm{x}^{\intercal}_{i}\widehat{\bm{\beta}}^{(n_{k-1})})-\dot{\Phi}(\bm{x}^{\intercal}_{i}{\bm{\beta}}^{*})\right\}\right].$
The uniformly boundedness of function $\ddot{\Phi}(t)$ implies that
$|\dot{\Phi}(\bm{x}^{\intercal}_{i}\widehat{\bm{\beta}}^{(n_{k-1})})-\dot{\Phi}(\bm{x}^{\intercal}_{i}{\bm{\beta}}^{*})|\leq M_{\Phi}\left|\bm{x}^{\intercal}_{i}(\widehat{\bm{\beta}}^{(n_{k-1})}-{\bm{\beta}}^{*})\right|\leq M_{\Phi}\lVert\bm{x}_{i}\lVert_{\infty}\lVert\widehat{\bm{\beta}}^{(n_{k-1})}-{\bm{\beta}}^{*}\lVert_{1}.$
Meanwhile,
$\left|\dot{\Phi}(\bm{x}^{\intercal}_{i}\widehat{\bm{\beta}}^{(n_{k-1})})+\dot{\Phi}(\bm{x}^{\intercal}_{i}{\bm{\beta}}^{*})-2y_{i}\right|\leq 2\left|\dot{\Phi}(\bm{x}^{\intercal}_{i}{\bm{\beta}}^{*})-y_{i}\right|+M_{\Phi}\lVert\bm{x}_{i}\lVert_{\infty}\lVert\widehat{\bm{\beta}}^{(n_{k-1})}-{\bm{\beta}}^{*}\lVert_{1},$
and
$\left|\bm{x}_{i}^{\intercal}{\widehat{\bm{\gamma}}_{\cdot,j}}^{(i)}\right|\leq\left|\bm{x}_{i}^{\intercal}\bm{\gamma}_{\cdot,j}^{*}\right|+\left\lVert\bm{x}_{i}\right\lVert_{\infty}\left\lVert{\widehat{\bm{\gamma}}_{\cdot,j}}^{(i)}-\bm{\gamma}_{\cdot,j}^{*}\right\lVert_{1}.$
Note that both $\bm{x}_{i}$, $\bm{x}_{i}^{\intercal}\bm{\gamma}_{\cdot,j}^{*}$ and $y_{i}$ are sub-gaussian. Applying Hoeffding’s inequality, we have
$\max_{i\in[n]}\lVert\bm{x}_{i}\lVert_{\infty}\leq c\sqrt{\log p},\ \max_{i\in[n]}\lVert\bm{x}_{i}^{\intercal}\bm{\gamma}_{\cdot,j}^{*}\lVert_{\infty}\leq c\sqrt{\log p},\ \max_{i\in[n]}|y_{i}-\mathbb{E}(y_{i})|\leq c\sqrt{\log p}$
holds with probability at least $1-p^{-4}$ for some universal constant $c$. Then the upper bound for (28) is given by
$\left\lvert\sum_{k=l+1}^{K}\sum_{i=n_{k-1}+1}^{n_{k}}\left[(\bm{x}_{i}^{\intercal}{\widehat{\bm{\gamma}}_{\cdot,j}}^{(i)})^{2}\left\{\dot{\Phi}(\bm{x}^{\intercal}_{i}\widehat{\bm{\beta}}^{(n_{k-1})})-y_{i}\right\}^{2}-(\bm{x}_{i}^{\intercal}{\widehat{\bm{\gamma}}_{\cdot,j}}^{(i)})^{2}\left\{\dot{\Phi}(\bm{x}^{\intercal}_{i}{\bm{\beta}}^{*})-y_{i}\right\}^{2}\right]\right\lvert$
$\leq$ $\sum_{k=l+1}^{K}\sum_{i=n_{k-1}+1}^{n_{k}}2\left|\bm{x}_{i}^{\intercal}{\widehat{\bm{\gamma}}_{\cdot,j}}^{(i)}\right|^{2}\left|\dot{\Phi}(\bm{x}^{\intercal}_{i}\widehat{\bm{\beta}}^{(n_{k-1})})+\dot{\Phi}(\bm{x}^{\intercal}_{i}{\bm{\beta}}^{*})-2y_{i}\right|\left|\dot{\Phi}(\bm{x}^{\intercal}_{i}\widehat{\bm{\beta}}^{(n_{k-1})})-\dot{\Phi}(\bm{x}^{\intercal}_{i}{\bm{\beta}}^{*})\right|$
$\lesssim$ $\sum_{k=l+1}^{K}\sum_{i=n_{k-1}+1}^{n_{k}}\Bigg{[}\left\{\sqrt{\log p}+(s_{1}d_{1}+s_{1}d_{0}+s_{0}d_{0})\frac{(\log p)^{2}}{\sqrt{n_{k-1}}}\right\}^{2}$
×$\ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \times\left\lvert\sqrt{\log p}+s_{0}d_{0}\frac{(\log p)^{2}}{\sqrt{n_{k-1}}}\right\lvert\times\left\lvert s_{0}d_{0}\frac{(\log p)^{2}}{\sqrt{n_{k-1}}}\right\lvert\Bigg{]}$
$\lesssim$ $\sum_{k=l+1}^{K}\Big{\{}s_{0}d_{0}\sqrt{(\log p)^{7}n_{k}}+s_{0}^{2}d_{0}^{2}(\log p)^{5}+s_{0}d_{0}(s_{1}d_{1}+s_{1}d_{0}+s_{0}d_{0})^{2}\frac{(\log p)^{6.5}}{\sqrt{n_{k}}}$
$\ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ +s_{0}^{2}d_{0}^{2}(s_{1}d_{1}+s_{1}d_{0}+s_{0}d_{0})^{2}\frac{(\log p)^{8}}{n_{k}}\Big{\}}$
$\lesssim$ $\ s_{0}d_{0}\sqrt{(\log p)^{7}n}+s_{0}^{2}d_{0}^{2}(\log p)^{5}\log_{2}(n)+s_{0}d_{0}(s_{1}d_{1}+s_{1}d_{0}+s_{0}d_{0})^{2}{(\log p)^{6.5}}$
$\ \ \ \ \ +s_{0}^{2}d_{0}^{2}(s_{1}d_{1}+s_{1}d_{0}+s_{0}d_{0})^{2}{(\log p)^{8}}=o(n).$
For another part, the proof is similar. First of all,
$\left\lvert\sum_{k=l+1}^{K}\sum_{i=n_{k-1}+1}^{n_{k}}\left[(\bm{x}_{i}^{\intercal}{\widehat{\bm{\gamma}}_{\cdot,j}}^{(i)})^{2}\left\{\dot{\Phi}(\bm{x}^{\intercal}_{i}{\bm{\beta}}^{*})-y_{i}\right\}^{2}-(\bm{x}_{i}^{\intercal}\bm{\gamma}_{\cdot,j}^{*})^{2}\left\{\dot{\Phi}(\bm{x}^{\intercal}_{i}{\bm{\beta}}^{*})-y_{i}\right\}^{2}\right]\right\lvert$
$\leq$ $\sum_{k=l+1}^{K}\sum_{i=n_{k-1}+1}^{n_{k}}\left\{\dot{\Phi}(\bm{x}^{\intercal}_{i}{\bm{\beta}}^{*})-y_{i}\right\}^{2}\left\{\bm{x}_{i}^{\intercal}({\widehat{\bm{\gamma}}_{\cdot,j}}^{(i)}-\bm{\gamma}_{\cdot,j}^{*})\right\}\left\{2\bm{x}_{i}^{\intercal}\bm{\gamma}_{\cdot,j}^{*}+\bm{x}_{i}^{\intercal}({\widehat{\bm{\gamma}}_{\cdot,j}}^{(i)}-\bm{\gamma}_{\cdot,j}^{*})\right\}$
$\lesssim$ $\sum_{k=l+1}^{K}\sum_{i=n_{k-1}+1}^{n_{k}}\bigg{[}(\log p)\times(s_{1}d_{1}+s_{1}d_{0}+s_{0}d_{0})\frac{(\log p)^{2}}{\sqrt{n_{k-1}}}$
×$\ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \times\left\{\sqrt{\log p}+(s_{1}d_{1}+s_{1}d_{0}+s_{0}d_{0})\frac{(\log p)^{2}}{\sqrt{n_{k-1}}}\right\}\bigg{]}$
$\lesssim$ $\sum_{k=l+1}^{K}\big{\{}(s_{1}d_{1}+s_{1}d_{0}+s_{0}d_{0})(\log p)^{3.5}\sqrt{n_{k}}+(s_{1}d_{1}+s_{1}d_{0}+s_{0}d_{0})^{2}(\log p)^{5}\big{\}}$
$\lesssim$ $(s_{1}d_{1}+s_{1}d_{0}+s_{0}d_{0})(\log p)^{3.5}\sqrt{n}+(s_{1}d_{1}+s_{1}d_{0}+s_{0}d_{0})^{2}(\log p)^{5}\log_{2}(n)=o(n).$
Combining the results for the two parts, we conclude (27). The proof of Theorem 3.3 is complete. ∎
Lemma C.1.
Under the conditions in Theorem 3.3, $|\Delta_{1}|=o_{\mathbb{P}}(\sqrt{n})$.
Lemma C.2.
Under the conditions in Theorem 3.3, $|\Delta_{2}|=o_{\mathbb{P}}(\sqrt{n})$.
Lemma C.3.
Under the conditions in Theorem 3.3, $|\Delta_{3}|=o_{\mathbb{P}}(\sqrt{n})$.
Lemma C.4.
Under the conditions in Theorem 3.3, $|\Delta_{4}|=o_{\mathbb{P}}(\sqrt{n})$.
The proofs of Lemmas C.1C.4 are deferred to the supplementary material.